OpenStudy (anonymous):

Find the antiderivative of the integral X^2(X^3+5)^9 dx

OpenStudy (anonymous):

\[\int\limits_{}^{} x ^{2}(x ^{3} + 5)^{9} dx \]

OpenStudy (anonymous):

let u=x^3+5

OpenStudy (anonymous):

so \[\int\limits_{}^{}x ^{2}(u)^{9}\] ??

OpenStudy (anonymous):

if u=x^3+5, then dx= (du/2x^2)

OpenStudy (anonymous):